# A quantum information theoretic analysis of reinforcement learning-assisted quantum architecture search

### v2

@@ Abstract
In the field of quantum computing, variational quantum algorithms (VQAs) represent a pivotal category of quantum solutions across a broad spectrum of applications. These algorithms demonstrate significant potential for realising quantum computational advantage. A fundamental aspect of VQAs involves formulating expressive and efficient quantum circuits (namely ansatz) and automating the search of such ansatz is known as quantum architecture search (QAS). RL-QAS involves optimising QAS using reinforcement learning techniques. This study investigates RL-QAS for crafting ansatzes tailored to the variational quantum state diagonalisation problem. Our investigation includes a comprehensive analysis of various dimensions, such as the entanglement thresholds of the resultant states, the impact of initial conditions on the performance of RL-agent, the phase change behaviour of correlation in concurrence bounds, and the discrete contributions of qubits in deducing eigenvalues through conditional entropy metrics. We leverage these insights to devise entanglement-guided admissible ansatz in QAS to diagonalise random quantum states using optimal resources. Furthermore, the methodologies presented herein offer a generalised framework for constructing reward functions within RL-QAS applicable to variational quantum algorithms.
@@ I Introduction
The ansatz design is critical, as it directly influences the expressivity and efficiency of the quantum solution. Given the vast potential configuration space of quantum circuits, the process of identifying a suitable ansatz is a challenging problem. In Ostaszewski et al. (2021); Kundu et al. (2024); Patel et al. (2024); Du et al. (2022), the search for the quantum circuits that solve various VQA has been automated using reinforcement learning (RL) techniques Sutton and Barto (2018). We particularly call ansatzes capable of solving a VQA an admissible ansatz. The automated search for such admissible ansatz using the RL-agent is termed reinforcement learning-assisted quantum architecture search (RL-QAS). In this setting, after a certain number of episodes, the RL agent returns possible configurations of quantum ansatz that solve the problem. For $E_{tot}$ episodes, the agent returns $E_{s}$ successful episodes, and in $E_{tot}-E_{s}$ episodes, the agent fails to find a configuration within a specific predefined accuracy and depth. It has been shown in Ostaszewski et al. (2021); Wu et al. (2023); Kundu et al. (2024); Patel et al. (2024); Kundu (2024) that classical machine learning-driven QAS algorithms outperform the state-of-art structures of ansatz and effectively optimise the number of parameters, depth, and impact of noise in VQAs.
@@
In this article, we investigate the admissible ansatzes proposed by the RL-agent from the perspective of quantum information theory Wilde (2013). For the investigation, we specifically focus on the reinforcement learning enhanced variational quantum state diagonalisation (RL-VQSD) problem Kundu et al. (2024). In the VQSD algorithm LaRose et al. (2019), the problem-inspired ansatz is based on density matrix exponentiation Lloyd et al. (2014), but its depth and number of gates increase exponentially with the number of qubits. Hence, searching for an optimal ansatz (with a minimum number of gates and depth) that diagonalises a quantum state is an open research problem. In RL-VQSD, the search for the optimal configuration of ansatz is done by using a double deep-Q network (DDQN). See Appendix B for further details on the VQSD algorithm, the agent and environment specifications. The RL-agent proposes $E_{s}<E_{tot}$ admissible ansatzes that solves the VQSD problem.
@@
Our investigation shows that the concurrence of the $E_{s}$ ansatz lies between an upper and lower bound, independent of the different initialisation of the double deep-Q network (DDQN), where each initialisation of DDQN corresponds to a different random initialisation of the weights. Moreover, we show that the upper and lower bounds are anti-correlated with respect to the increment in the initial state entanglement. The anti-correlation turns into a mild, and later strong, correlation when the initial state entanglement goes beyond $0.322$, indicating a phase change in correlation between the upper and lower bounds of the concurrence. The phase transition provides us with an in-depth insight into the relation between the input and the output of the VQSD algorithm. Further, our investigation reveals that the RL-agent can generate high concurrence admissible ansatz with fewer 2-qubit gates and circuit depth compared to ansatzes with very low concurrence. This is used as a guiding principle to design a two-part ansatz, where an entanglement enhancing initial block adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$. Furthermore, we quantify the contribution of individual qubits of the RL-ansatz using conditional quantum entropy. Using such a measure, we observe a correlation between entanglement and quantum entropy, revealing why RL-VQSD can efficiently find the largest eigenvalues but fails to find the smallest ones.
@@
The structure of the paper is as follows. We present an overview of our main results in Sec. II. In Sec. III.1, we analyse the upper and lower bounds on the concurrence of the admissible ansatzes that solve the VQSD problem. Specifically, in Sec. III.1.1 we present the variation of the entanglement bounds for different weight initialisation of the DDQN. We observe the correlation properties of the entanglement bounds dependent on the concurrence of the initial state in Sec. III.1.2. In Sec. III.1.3, we present the dependence of the entanglement bounds on the number of gates and circuit depth of the ansatz. The analysis of the contribution of individual qubits to the admissible ansatz is provided in Sec. III.2. In particular, we present the contribution of each qubit for change in the entanglement of the ansatz in Sec. III.2.1. In Sec. III.2.2, we explain why RL-VQSD can efficiently find the largest eigenvalues of the initial state but not the smallest eigenvalues. We provide concluding remarks and discuss possible future directions in Sec. IV.
@@ II Contributions
    Demonstrating that the entanglement of the states generated by the RL-agent-proposed admissible ansatzes lies within a lower and upper concurrence bound. The upper and lower bounds of concurrence remain consistent across various weight initialisation of the DDQN. Furthermore, we show that these bounds are anti-correlated with respect to increasing entanglement until concurrence surpasses $0.322$; after this point, the anti-correlation turns into correlation. Hence, we introduce the concurrence $0.322$ as the phase change point.
@@
    We evaluate individual qubits’ contribution in diagonalising 2-qubit quantum states using conditional quantum entropy. Our investigation shows that the qubits have equal contributions on average in diagonalising random quantum states.
@@
    We further investigate the correlation in conditional entropy between qubits for different inferred eigenvalues. It requires mild anti-correlation for the first two largest eigenvalues and strong anti-correlation for the smallest eigenvalues. Most admissible ansatzes lie in a mild anti-correlation regime, explaining VQSD’s limitation in finding the smallest eigenvalues.
III Main results
In this section, we investigate the upper and lower bounds of concurrence of the ansatzes produced by the RL agent in diagonalising $2$-qubit random quantum states sampled from IBM qiskit’s random_density_matrix module. For the simulation, we consider multiple such quantum states (specified in the caption of figures), and for each state, we use $10000$ episodes of RL-VQSD to diagonalise the state. The configuration of the RL-VQSD is provided in detail in Appendix B.
III.1 Entanglement
III.1.1 DDQN weight independent entanglement bound
To benchmark the performance of the RL-agent for a specific problem, it is ideal to take multiple initialisations of the DDQN (see Appendix B for a brief description).
@@ III.1.1 DDQN weight independent entanglement bound
To efficiently investigate the upper and lower bounds of concurrence Hill and Wootters (1997); Wootters (1998) of the admissible ansatz, we calculate the concurrence of admissible ansatz, which is defined by the entanglement induced by the admissible ansatz when initialised with |$\ket{00}$. We first sample different random quantum states, then diagonalise them using RL-VQSD Kundu et al. (2024) for five different initialisations of DDQN weights. Thereafter, the agent suggests different possible configurations for the RL-ansatz, which diagonalises the sampled random quantum states. For a specific state, we collect such valid quantum circuits and find the concurrence of the state obtained after evolving through the RL-ansatz. It should be noted that finding the concurrence of the evolved state indicates how much the RL-ansatz induces entanglement to the input sampled state. See Appendix C for further analysis of the sampled random quantum states, where we show that in qiskit, random_density_matrix even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low.
@@
III.1.2 Anti-correlation between entanglement bounds
@@
In Figure 1, we illustrate the variation of the entanglement bounds with the entanglement of the input quantum state. To investigate the dependency, we use the Pearson correlation coefficient (PCC) Lee Rodgers and Nicewander (1988) as a quantifier of correlation. See Appendix A for a brief description. We see that the upper and lower bounds of entanglement are initially strongly anti-correlation when the input state entanglement is below $0.322$ and then strongly correlate afterwards, indicating a phase change in the upper and lower bounds of entanglement.
@@
We present in Figure 2, the variation of $\eta_{0}^{k}$ (in blue) and $\eta_{1}^{k}$ (in red) for different values of $k$. We observe that $\eta_{1}^{k}\approx-1~{}\forall k\in(i,j)$. This indicates strong cumulative anti-correlation between the upper and lower bounds of concurrence. For $\eta_{0}^{k}$, we observe a gradual change of $\text{PCC}_{kj}$ from -0.582 to 1. We note that $\eta_{0}^{k}$ changes from $-0.0451$ to $0.279$ for $k\in[0.313,0.379]$ with $\eta_{0}^{k}=0$ at $k=k_{\ast}=0.322$. We call this point the phase transition point of correlation between upper and lower bounds of concurrence for the VQSD ansatz. At values of $k<k_{\ast}$, there is anti-correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$. While for values of $k>k_{\ast}$, there is (anti-)correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $(\mathbbm{u}_{2})~{}\mathbbm{u}_{1}$.
III.1.3 Enhancing the performance of agent based on initial state entanglement
@@
Figure 3: The RL-agent can generate an ansatz with strong entanglement with fewer 2-qubit gates and circuit depth compared to the ansatzes with weak entanglement. As the main aim of RL-VQSD is to diagonalise unitary with a small number of gates and depth, our analysis shows that if we particularly focus on admissible ansatzes near the upper bound of concurrence, we can not only reduce the computational cost of RL-VQSD severely but find optimal admissible ansatzes. The numbers in the bars correspond to the amount of entanglement generated by the RL-ansatz when initialised in the vacuum state.
Our investigation reveals that the admissible ansatzes corresponding to the upper bound of concurrence has, on average, a shorter depth of $13.23$ and a smaller number of gates of $18.5$ compared to the admissible ansatzes corresponding to the lower bound of concurrence where the depth is $13.9$ and number of gates is $19$ over $25$ random quantum states. Moreover, Figure 3 makes it more prominent that the minimum number of gates and the depth of the admissible ansatzes are smaller for the upper bound of concurrence than the lower. This observation is crucial in reducing the subspace of all possible admissible ansatzes. The main motivation of RL-QAS in VQAs is to find ansatzes that solve a VQA with a minimum number of gates and circuit depth; in the case of RL-VQSD, this observation can help us to narrow down the search for an optimal admissible ansatz in the two following ways.
@@
In RL-VQSD Kundu et al. (2024), the RL-state that encodes the ansatz initiates from an empty circuit (i.e., in circuit 2 $U_{\text{RL-agent}}=\mathds{1}$). We showed that the optimal admissible ansatz, on average, lies in the upper bound of concurrence, motivating us to initialise the quantum state (that is diagonalised using RL-VQSD) closest to the upper bound of entanglement. To do this, the ansatz is divided into two parts
@@
where the first, Entanglement Enhancing (EE) block, helps the diagonalising state to achieve the highest entanglement before it is fed to the second part, $U_{\text{RL-agent}}$, which is decided by the RL-agent. For example, for seed $27$ (see Appendix C for details), we get a random quantum state with concurrence $0.198$, and using the EE, we can increase the input state concurrence to $0.215$ at $\theta=0.5$ as shown in Fig. 4.
@@
Table 2: Turning on the Entanglement Enhancing (EE) mode (denoted as Enhanced ($\theta=0.5$)) enhances the learning of the performance of the RL-agent in the absence of the EE mode (denoted as Default ($\theta=0$)). In the first $3500$, the Enhanced ($\theta=0.5$) collects $2.08$ times more reward (tot. rwd. coll.) and $2.22$ times more successful episodes (Tot. succ. ep.) than Default ($\theta=0$). Moreover, the average number of 2-qubit (2q gate) and depth over the first $100$ successful episodes reduced in the case of Enhanced ($\theta=0.5$) settings compared to Default ($\theta=0$), showing enhanced learning of the RL-agent for entanglement enhanced input state.
@@
III.2 Conditional quantum entropy
III.2.1 Contribution of individual qubits for change in entanglement
We quantify the contribution of each qubit using the conditional quantum entropy $S_{q_{0}|q_{1}}^{\rho^{\prime}_{q_{0}q_{1}}}$ Cerf and Adami (1997, 1999); Wilde (2013). The conditional entropy of an individual qubit for an $N$ qubit quantum system indicates the amount of uncertainty remaining about that qubit after measuring the whole system. Hence, it is essential to know how much information the state of qubits provides about the state of the $N$ qubits. For a detailed discussion and notation, see Appendix A.
Figure 5: Relative change in the contribution of qubits 0 and 1 as a function of change in entanglement. We quantify the relative contribution of individual qubits to the RL-VQSD ansatz using $\Delta_{q_{0}}^{q_{1}}$ (see Eq. (3)). For the change in concurrence due to the ansatz, the distribution of $\Delta_{q_{0}}^{q_{1}}$ is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$. The magnitude of the cumulative weight of all data points in the region $\Delta_{q_{0}}^{q_{1}}<0$ is $4.854$ while that for data points in the region $\Delta_{q_{0}}^{q_{1}}>0$ is $5.524$. This indicates equal contribution from both the qubits to the RL-VQSD ansatz for the range of change in entanglement.

### v3

@@ Abstract
In the field of quantum computing, variational quantum algorithms (VQAs) represent a pivotal category of quantum solutions across a broad spectrum of applications. These algorithms demonstrate significant potential for realising quantum computational advantage. A fundamental aspect of VQAs involves formulating expressive and efficient quantum circuits (namely ansatz), and automating the search of such ansatz is known as quantum architecture search (QAS). Recently, reinforcement learning (RL) techniques is utilized to automate the search for ansatzes, known as RL-QAS. This study investigates RL-QAS for crafting ansatz tailored to the variational quantum state diagonalisation problem. Our investigation includes a comprehensive analysis of various dimensions, such as the entanglement thresholds of the resultant states, the impact of initial conditions on the performance of RL-agent, the phase transition behaviour of correlation in concurrence bounds, and the discrete contributions of qubits in deducing eigenvalues through conditional entropy metrics. We leverage these insights to devise an entanglement-guided admissible ansatz in QAS to diagonalise random quantum states using optimal resources. Furthermore, the methodologies presented herein offer a generalised framework for constructing reward functions within RL-QAS applicable to variational quantum algorithms.
@@ I Introduction
The ansatz design is critical, as it directly influences the expressivity and efficiency of the quantum solution. Given the vast potential configuration space of quantum circuits, the process of identifying a suitable ansatz is a challenging problem. In Ostaszewski et al. (2021); Kundu et al. (2024); Patel et al. (2024); Du et al. (2022), the search for the quantum circuits that solve various VQA has been automated using reinforcement learning (RL) techniques Sutton and Barto (2018). We particularly call ansatzes capable of solving a VQA an admissible ansatz. The automated search for such an admissible ansatz using the RL-agent is termed reinforcement learning-assisted quantum architecture search (RL-QAS). In this setting, after a certain number of episodes, the RL agent returns possible configurations of quantum ansatz that solve the problem. For $E_{tot}$ episodes, the agent returns $E_{s}$ successful episodes, and in $E_{tot}-E_{s}$ episodes, the agent fails to find a configuration within a specific predefined accuracy and depth. It has been shown in Ostaszewski et al. (2021); Wu et al. (2023); Kundu et al. (2024); Patel et al. (2024); Kundu (2024) that classical machine learning-driven QAS algorithms outperform the state-of-art structures of ansatz and effectively optimise the number of parameters, depth, and impact of noise in VQAs.
@@
In this article, we numerically investigate the admissible ansatzes proposed by the RL-agent from the perspective of quantum information theory Wilde (2013). For the investigation, we specifically focus on the reinforcement learning enhanced variational quantum state diagonalisation (RL-VQSD) problem Kundu et al. (2024). In the VQSD algorithm LaRose et al. (2019), the problem-inspired ansatz is based on density matrix exponentiation Lloyd et al. (2014), but its depth and number of gates increase exponentially with the number of qubits. Hence, searching for an optimal ansatz (with a minimum number of gates and depth) that diagonalises a quantum state is an open research problem. In RL-VQSD, the search for the optimal configuration of ansatz is done by using a double deep-Q network (DDQN). See Section III for further details on the VQSD algorithm, the agent and environment specifications. The RL-agent proposes $E_{s}<E_{tot}$ admissible ansatzes that solves the VQSD problem. The primary motivation of our work is to filter potential admissible ansatzes based on concurrence, using the highest and lowest entanglement bounds. This helps identify the range of optimal ansatzes and determine if increasing entanglement can improve the agent’s performance. Another goal is to see if we can find an admissible ansatz that both diagonalizes and enhances the entanglement of the vacuum state.
@@
Our investigation shows that the concurrence of the $E_{s}$ ansatz lies between an upper and lower bound, independent of the different initialisation of the double deep-Q network (DDQN), where each initialisation of DDQN corresponds to a different random initialisation of the weights. Moreover, through the numerical analysis we show that the upper and lower bounds are anti-correlated with respect to the increment in the initial state entanglement. The anti-correlation turns into a mild, and later strong, correlation when the initial state entanglement goes beyond $0.322$, indicating a phase transition in correlation between the upper and lower bounds of the concurrence. The phase transition provides us with an in-depth insight into the relation between the input and the output of the VQSD algorithm. Further, our investigation reveals that the RL-agent can generate high concurrence admissible ansatzes with fewer 2-qubit gates and circuit depth compared to ansatzes with very low concurrence. This is used as a guiding principle to design a two-part ansatz, where an entanglement enhancing initial block adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$. Furthermore, we quantify the contribution of individual qubits of the RL-ansatz using conditional quantum entropy. Using such a measure, we observe a correlation between entanglement and quantum entropy, revealing why RL-VQSD can efficiently find the largest eigenvalues but fails to find the smallest ones.
@@
The structure of the paper is as follows. We present an overview of our main results in Sec. II. In Sec. III, we present an overview of the reinforcement learning enhanced variational quantum state diagonalisation algorithm. We present an analysis of the sampled 2-qubit Haar-random quantum states in Sec. IV. We present the main results in Sec. V. Specifically, in Sec. V.1, we analyse the upper and lower bounds on the concurrence of the admissible ansatzes that solve the VQSD problem. In Sec. V.1.1 we present the variation of the entanglement bounds for different weight initialisation of the DDQN. We observe the correlation properties of the entanglement bounds dependent on the concurrence of the initial state in Sec. V.1.2. In Sec. V.1.3, we present the dependence of the entanglement bounds on the number of gates and circuit depth of the ansatz. The analysis of the contribution of individual qubits to the admissible ansatz is provided in Sec. V.2. In particular, we present the contribution of each qubit for change in the entanglement of the ansatz in Sec. V.2.1. In Sec. V.2.2, we explain why RL-VQSD can efficiently find the largest eigenvalues of the initial state but not the smallest eigenvalues. We provide concluding remarks and discuss possible future directions in Sec. VI.
@@ II Contributions
    We numerically demonstrate that the entanglement of the states generated by the RL-agent-proposed admissible ansatzes lies within a lower and upper concurrence bound. The upper and lower bounds of concurrence remain consistent across various weight initialisation of the DDQN. Furthermore, we show that these bounds are anti-correlated with respect to increasing entanglement until concurrence surpasses $0.322$; after this point, the anti-correlation turns into correlation. Hence, we introduce the concurrence $0.322$ as the phase transition point.
@@
    We numerically evaluate individual qubits’ contribution in diagonalising 2-qubit quantum states using conditional quantum entropy. Our investigation shows that the qubits have equal contributions on average in diagonalising random quantum states.
@@
    We further investigate the correlation in conditional entropy between qubits for different inferred eigenvalues. It requires mild correlation for the first two largest eigenvalues and mild anti-correlation for the smallest eigenvalues. Most admissible ansatzes lie in a mild correlation regime, explaining VQSD’s limitation in finding the smallest eigenvalues.
III The reinforcement learning enhanced variational quantum state diagonalisation
where $\rho_{\text{diag}}$ is $\rho$ diagonalised in its eigenbasis and $\vec{\theta}_{\text{opt}}$ are the optimal angles. Classical gradient-based methods such as SPSA and Gradient Descent, or gradient-free optimisation Powell (1994, 2006) can be utilised in the training process. (2) In the eigenvalue readout subroutine, using the optimised unitary $U(\vec{\theta}_{\text{opt}})$ and one copy of state $\rho$, one can extract all the eigenvalues for low-rank states or the largest eigenvalues for full-rank states. This is achieved by measuring $\rho^{\prime}$ in the computational basis, $\mathbf{b}=b_{1}b_{2}\ldots b_{n}$, as follows
$\lambda^{\prime}=\bra{\mathbf{b}}\rho^{\prime}\ket{\mathbf{b}},$
The key update equation is
IV Analysis of the Haar random input quantum states
In this section, we analyse the 2-qubit Haar random quantum states. The Haar random states can be efficiently simulated using polynomial-time algorithms, as demonstrated in prior research Alagic et al. (2020). This holds even in scenarios where adversaries have unbounded capabilities with black-box oracle access Chen et al. (2024). Here, we sample $100000$ Haar random quantum states and observe how the eigenvalues and the qubits’ conditional entropy change with the state’s entanglement.
In Fig. 1, see that the difference in conditional entropy in qubits of the input state varies evenly on either side of $S_{AB}^{\rho_{AB}}=0$ where $A$ and $B$ are subsystems containing one qubit each and $\rho_{AB}$ is the sampled Haar random state. Meanwhile, the area of the distribution of $S_{AB}^{\rho_{AB}}$ on either side of zero entropy converges towards zero as the input state entanglement increases. Furthermore, the magnitude of the input quantum state’s largest eigenvalue increases with the increase in entanglement, and all the other eigenvalues converge towards 00. It also shows that even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low, justifying the low density of points after $0.6$ in Fig. 2.
V Main results
In this section, we investigate the upper and lower bounds of concurrence of the ansatzes produced by the RL agent in diagonalising $2$-qubit random quantum states sampled from IBM qiskit’s random_density_matrix module. For the simulation, we consider multiple such quantum states (specified in the caption of figures), and for each state, we use $10000$ episodes of RL-VQSD algorithm to diagonalise the state. The algorithm consists of DDQN with $\epsilon$-greedy policy where the $\epsilon$ decays in each step by a factor of $0.99995$ from its initial value $1$, down to a minimal value $0.05$. Furthermore, the network is trained by ADAM optimizer Kingma and Ba (2014) with a learning rate of 0.0001. The neural network consists of 5 hidden layers with 1000 neurons and a buffer size of 15000. The configuration of the RL-VQSD is provided in detail in Section III.
V.1 Entanglement
V.1.1 DDQN weight independent entanglement bound
To benchmark the performance of the RL-agent for a specific problem, it is ideal to take multiple initialisations of the DDQN (see Section III for a brief description).
@@ V.1.1 DDQN weight independent entanglement bound
To efficiently investigate the upper and lower bounds of concurrence Hill and Wootters (1997); Wootters (1998) of the admissible ansatz, we calculate the concurrence of the admissible ansatz, which is defined by the entanglement induced by the admissible ansatz when initialised with |$\ket{00}$. We first sample different random quantum states, then diagonalise them using RL-VQSD Kundu et al. (2024) for five different initialisations of DDQN weights. Thereafter, the agent suggests different possible configurations for the RL-ansatz, which diagonalises the sampled random quantum states. For a specific state, we collect such valid quantum circuits and find the concurrence of the state obtained after evolving through the RL-ansatz. It should be noted that finding the concurrence of the evolved state indicates how much the RL-ansatz induces entanglement to the input sampled state. See Section IV for further analysis of the sampled random quantum states, where we show that in qiskit, random_density_matrix even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low.
@@
V.1.2 Anti-correlation between entanglement bounds
@@
In Fig. 2, we illustrate the variation of the entanglement bounds with the entanglement of the input quantum state. To investigate the dependency, we use the Pearson correlation coefficient (PCC) Lee Rodgers and Nicewander (1988) as a quantifier of correlation. See Section A for a brief description. We see that the upper and lower bounds of entanglement are initially strongly anti-correlation when the input state entanglement is below $0.322$ and then strongly correlate afterwards, indicating a phase transition in the upper and lower bounds of entanglement.
@@
Intuitively, when the entanglement of the starting state is low i.e. $k\leq 0.322$, it becomes easier for the dephasing subroutine of the VQSD algorithm to eliminate the off-diagonal terms. Hence the agent proposes an admissible ansatz to either convert the state into a separable state (eliminating the need for a dephasing gate at the end) or increase its entanglement to the maximum level, thereby making full use of the dephasing gate’s potential. This dynamic explains the observed anti-correlation between the entanglement bounds.
As noted previously, when the entanglement of the starting state is high (in the region $k\geq 0.322$), it becomes easier for the dephasing subroutine to eliminate the off-diagonal terms. Thus for states with high entanglement, the agent proposes an admissible ansatz to maximize the entanglement of the starting state. Hence we observe the positive relation between the entanglement bounds.
@@
We present in Fig. 3, the variation of $\eta_{0}^{k}$ (in blue) and $\eta_{1}^{k}$ (in red) for different values of $k$. We observe that $\eta_{1}^{k}\approx-1~{}\forall k\in(i,j)$. This indicates strong cumulative anti-correlation between the upper and lower bounds of concurrence. For $\eta_{0}^{k}$, we observe a gradual change of $\text{PCC}_{kj}$ from -0.582 to 1. We note that $\eta_{0}^{k}$ changes from $-0.0451$ to $0.279$ for $k\in[0.313,0.379]$ with $\eta_{0}^{k}=0$ at $k=k_{\ast}=0.322$. We call this point the phase transition point of correlation between upper and lower bounds of concurrence for the VQSD ansatz. At values of $k<k_{\ast}$, there is anti-correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$. While for values of $k>k_{\ast}$, there is (anti-)correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $(\mathbbm{u}_{2})~{}\mathbbm{u}_{1}$.
The investigation in Fig. 2 and Fig. 3 suggests that the structure of the quantum state in terms of its entanglement properties can significantly influence the effectiveness of certain ansatz strategies. Specifically, when the entanglement is low i.e. below $0.322$, the variability in the entanglement bounds can be exploited. For wide entanglement bounds, the flexibility of the ansatz allows for the generation of states that are either highly entangled or nearly separable, starting from $|0\rangle^{\otimes 2}$.
V.1.3 Enhancing the performance of agent based on initial state entanglement
@@
Figure 4: The RL-agent can generate an admissible ansatz with strong entanglement with fewer 2-qubit gates (2-qubit gate) and circuit depth (Depth) and total number of gates (Numb. gate) compared to the ansatzes with weak entanglement. As the main aim of RL-VQSD is to diagonalise unitary with a small number of gates and depth, our analysis shows that if we particularly focus on the admissible ansatzes near the upper bound of concurrence, we can not only reduce the computational cost of RL-VQSD severely but also find the optimal admissible ansatzes.
Our investigation reveals that the admissible ansatzes corresponding to the upper bound of concurrence has, on average, a shorter depth of $13.23$ and a smaller number of gates of $18.5$ compared to the admissible ansatzes corresponding to the lower bound of concurrence where the depth is $13.9$ and number of gates is $19$ over $25$ random quantum states. Moreover, Fig. 4 makes it more prominent that the minimum number of gates and the depth of the admissible ansatzes are smaller for the upper bound of concurrence than the lower. This observation is crucial in reducing the subspace of all possible admissible ansatzes. The main motivation of RL-QAS in VQAs is to find ansatzes that solve a VQA with a minimum number of gates and circuit depth; in the case of RL-VQSD, this observation can help us to narrow down the search for an optimal admissible ansatz in the two following ways.
@@
In RL-VQSD Kundu et al. (2024), the RL-state that encodes the ansatz initiates from an empty circuit (i.e., in circuit 5 the Entanglement Enhancing block is absent and $U_{\text{RL-agent}}=\mathds{1}$). We showed that the optimal admissible ansatz, on average, lies in the upper bound of concurrence, motivating us to initialise the quantum state (that is diagonalised using RL-VQSD) closest to the upper bound of entanglement. To do this, the ansatz is divided into two parts
@@
where the first, Entanglement Enhancing (EE) block, helps the diagonalising state to achieve the highest entanglement before it is fed to the second part, $U_{\text{RL-agent}}$, which the RL-agent decides. For example, for seed $27$ (see Section IV for details), we get a random quantum state with concurrence $0.198$ (for seed $27$, and see Appendix A.4 for details on how to generate Haar random state), and using the EE, we can increase the input state concurrence to $0.215$ at $\theta=0.5$ as shown in Fig. 5.
@@
Table 2: Turning on the Entanglement Enhancing (EE) mode (denoted as Enhanced ($\theta=0.5$)) enhances the learning of the performance of the RL-agent in the absence of the EE mode (denoted as Default ($\theta=0$)). In the first $3500$ episodes, the Enhanced ($\theta=0.5$) collects $2.08$ times more reward (tot. rwd. coll.) and $2.22$ times more successful episodes (Tot. succ. ep.) than Default ($\theta=0$). Moreover, the average number of 2-qubit (2q gate) and depth over the first $100$ successful episodes reduced in the case of Enhanced ($\theta=0.5$) settings compared to Default ($\theta=0$), showing enhanced learning of the RL-agent for entanglement enhanced input state.
@@
It should be noted that due to the sake of the interpretability of the EE block, we chose the block resembling a maximally entangled circuit for a 2-qubit system. Where the entangler (i.e. the CX) gate is parameterised and can induce no entanglement (for $\theta=0$) and maximal entanglement (for $\theta=0.5$) to $|00\rangle$ state. The second Hadamard (H) gate on the first qubit is introduced to cancel out the first H in the absence of the entangler at $\theta=0$. We can automate the search for gateset in the EE block utilising an RL-agent and diagonalise the input state simultaneously by decomposing the RLQAS into two subroutines. Suppose the total number of steps in an episode is $S$, then in the first subroutine for $s_{\text{EE}}<S$ steps the agent’s task would be to find a gateset (consists of maximum $s_{\text{EE}}$ gates) that maximises the entanglement of the input state by utilising a reward function $R_{\text{EE}}$. One can formulate $R_{\text{EE}}$ similar to ref. Kuo et al. (2021). Wherein in the second subroutine, for the remaining steps $S-s_{\text{EE}}$, the agent focuses on finding possible structures of $U_{\text{RL-ansatz}}$ to diagonalise the output state of the EE block using the reward defined in Eq. 2.
@@
V.2 Conditional quantum entropy
V.2.1 Contribution of individual qubits for change in entanglement
We quantify the contribution of each qubit using the conditional quantum entropy $S_{q_{0}|q_{1}}^{\rho^{\prime}_{q_{0}q_{1}}}$ Cerf and Adami (1997, 1999); Wilde (2013). The conditional entropy of an individual qubit for an $N$ qubit quantum system indicates the amount of uncertainty remaining about that qubit after measuring the whole system. Hence, it is essential to know how much information the state of qubits provides about the state of the $N$ qubits. For a detailed discussion and notation, see Section A.
Figure 6: We analyze the relative change in the contribution of the first and second qubits as a function of changes in entanglement. To quantify the relative contribution of individual qubits to the RL-VQSD ansatz, we use $\Delta_{q_{0}}^{q_{1}}$ (see Eq. (6)). For the change in concurrence due to the ansatz, the distribution of $\Delta_{q_{0}}^{q_{1}}$ is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$. The cumulative weight of all data points in the region $\Delta_{q_{0}}^{q_{1}}<0$ is 4.854, while that for data points in the region $\Delta_{q_{0}}^{q_{1}}>0$ is 5.524. This indicates an equal contribution from both qubits to the RL-VQSD ansatz across the range of changes in entanglement.